Layout instructions (TEX) for the Proceedings of the ECM Budapest 1996

The purpose of the following simple instructions is to guarantee a minimum of standardization and to make it easier for us to work with your TEX-file. However, we do not intend to bother the authors with details of the final layout. It will be our job to obtain a unified style by adapting the contributions of all speakers. If you use LATEX or AMSTEX you are welcome to work with the standard styles “article” and “amsppt”, respectively. Instead of providing you with an additional complicated macro package we will use modifications of these styles to achieve a new layout.

General remarks. Please include no extra macros in your file except those which you are actually going to use. Moreover, please make sure not to have several macros for the same thing but stay with one of them.

We gratefully appreciate if the text does not contain “overfull” and/or “underfull” boxes, if equations do not exceed the indicated width, if hyphenations have been checked, and if the hierarchical structure of the book is clear. There is no need for marking each level with different fonts or special font sizes. Please avoid caps and underlines and do not use \hoffset or\voffset, nor true sizes such as, for example, truecm.

General layout. Plain TEX: 10pt (default-)fonts

\hsize=120mm \vsize=185mm \parindent=8mm \frenchspacing

equations centered, equation numbers on the right side (default in plain TEX).

AMSTEX: \documentstyle{amsppt} \pagewidth{120mm} \pageheight{185mm} \parindent=8mm \frenchspacing \tenpoint

equations centered, equation numbers (tags) on the right side (default in AMSTEX).

LATEX: \documentstyle{article} (10pt, default) \setlength{\textwidth}{120mm} \setlength{\textheight}{185mm} \setlength{\parindent}{8mm} \setlength{\oddsidemargin}{0pt} \frenchspacing \pagestyle{plain}

equations centered, equation numbers on the right side (default in article-style). If you introduce labels for cross-referencing please make sure to use them accordingly rather than to insert the numbers that are referred to by hand.

Page numbering, running heads, footnotes. Since we will add running heads later you can simply leave the headers blank. Just use the standard page numbering. It would be helpful though, if you added in a comment an abbreviated version of the title which we can use for the running heads. For footnotes, please choose the same font as for the main text, but in a smaller size.

References. Here are some models.

[1] Collected Works of Raul Bott. Volume 1: Topology and Lie Groups, (R. MacPherson, ed.), Birkhäuser Boston, Basel, Berlin 1993.

[2] R. Bott, Stable Homotopy of the Classical Groups, Ann. of Math., 70 (1959), 313–337.

⋯

[B] R. Bott, An Application of the Morse Theory to the Topology of Lie Groups, Proceedings of the International Congress of Mathematicians Edinburgh 1958, (1958), 14–21.

[N] E. Noether, Hyperkomplexe Systeme in Ihren Beziehungen zur kommutativen Algebra und Zahlentheorie, Verhandlungen des Internationalen Mathematiker-Kon-gresses Zürich 1932. I. Band, (1933), 189–194.

Indentation. Though indentation to indicate a new paragraph is welcome, please do not use indentation when the new paragraph is already marked by an extra vertical space, as for example in the case of the first paragraph following a heading. Unless the style-file you are using is taking care of this already, please insert \noindent in such a situation. Likewise, please make sure not to indent the first line of theorems, lemmas, proofs etc., but rather to separate them from the preceding text by a little extra vertical space. Beware of inserting an empty line before or after a display in your source file. This has an effect on indentation as well as on TEX’s choice of pagebreaks.

Figures and tables. Please send figures as

(a) fair copy on paper, if possible scaled to about 200%, or

(b) as EPS-file (encapsulated postscript).

It is sometimes difficult to place insertions at an exact location in the final form of the book. Therefore, all figures and tables should be numbered and you should refer to these numbers within the text. Please avoid formulations like “the following figure … ”.

Transmission. Please send us a print-out of your work on paper, and all the TEX-files including all .tex, .dvi, .log, .aux files as well as your own personal macros and input files which are not commonly known (e.g. article.sty is commonly known). We need the .log files since they contain relevant information on the used software (e.g. which version of LATEX has been used). You can send the files either electronically or on floppy disk(s) in DOS or MacIntosh-format.

For further information please contact:  Birkhäuser Verlag AG  Tel.: ++41/61/205 07 07
Dr. Thomas Hintermann  P.O. Box 133  Fax: ++41/61/205 07 99
Editorial Department  CH-4010 Basel/Switzerland  e-mail: hintermann@birkhauser.ch